OpenStudy (anonymous):

I REALLY need help with Equations with Rational Expressions: Comment if you know how to do them and then i will put up the question! Thanks!

OpenStudy (anonymous):

\[\frac{ 5 }{ 2y+6 }+\frac{ 1 }{ y-2 }=\frac{ 3 }{ y+3 }\]

OpenStudy (anonymous):

WHAT'S THE QUESTION?

OpenStudy (anonymous):

Its in the Comments, did it not show up?